Money demand in an economy in which no interest is


paid on money is


Md


p == 500 + 0.2Y - 1000i.


a. Suppose that P = 100, Y = 1000, and i = 0.10. Find


real money demand, nominal money demand, and


velocity.


b. The price level doubles from P = 100 to P = 200.


Find real money demand, nominal money


demand, and velocity.

1
Expert's answer
2021-11-19T10:30:35-0500

Given:

The money demand is:"p = 500 + 0.2Y \u2212 1,000i"

Now, the P is = 100

Y = 1,000

I = 0.10

Part-A)

  • The real money demand, nominal money demand, and velocity:

"Real Money Demand = \frac{MD}{P}\\ = 500 + 0.2Y − 1,000i\\ = 500 + 0.2 (1,000) − 1,000 (0.10)\\ = 500 + 200 − 100\\ = 700 − 100\\ = 600"


"Nominal\space Money\space Demand = Real\space money \space demand × Price\\ = 600 × 100\\= 60,000"


"Velocity:\\

MV = PY\\60,000 × V = 100 × 1,000\\60,000 × V = 100,000\\V = \frac{100,000}{60,000}\\V = 1.66"


Part-B)

  • The price level doubles from P = 100 to P = 200, then the real money demand, nominal money demand, and velocity:

"Real Money Demand = \frac{MD}{P}\\ = 500 + 0.2Y − 1,000i\\ = 500 + 0.2 (1,000) − 1,000 (0.10)\\ = 500 + 200 − 100\\ = 700 − 100\\ = 600"

"Nominal\space Money\space Demand = Real\space money \space demand × Price\\ = 600 × 200\\= 120,000"


"Velocity:\\

MV = PY\\120,000 × V = 200 × 1,000\\120,000 × V = 200,000\\V = \frac{200,000}{120,000}\\V = 1.66"